What is the 5th term of the following arithmetic sequence: -3x-8, 2x-1, 7x+6
Answer by stanbon(75887)

What is the 5th term of the following arithmetic sequence: -3x-8, 2x-1, 7x+6
----
a(1) = -3x-8
d = (2x-1)-(-3x-8) = 5x+7
-----
Ans: a(5) = a(1)+4d
a(5) = -3x-8+4(5x+7) = -3x+20x -1 = 17x-1
--------------------------
Cheers,
Stan H.
-------------